### v1

# STUPP-24-271

Abstract
In this letter, we perform an almost general analysis of flavor mixing matrices to investigate the discriminative power of CP phases by the next-generation neutrino oscillation experiments. As an assumption, we neglect the 1-3 mixing of diagonalization for more hierarchical fermions $u,e$. Thus there are two sources of CP violation in the flavor mixing matrices, the intrinsic CP phase $\delta_{d,\nu}$ in diagonalization of less hierarchical fermions $d,\nu$ and relative phases between two unitary matrices.
By eliminating unphysical phases and imposing constraints of the three measured mixing angles, the flavor mixing matrices are analytically displayed by two phases and the 1-2 mixing $s_{u,e}$ of more hierarchical fermions. For sufficiently small 1-2 mixing $s_{e}$ of charged leptons, the Dirac phase $\delta$ is almost identical to the intrinsic phase of neutrinos $\delta_{\nu}$. Therefore, future detection of the Dirac phase by the Hyper-Kamiokande experiment indicates the observation of $\delta_{\nu}$. On the other hand, if such a CP violation is not observed, an upper limit is placed on a combination of $\delta_{\nu}$ and relative phases.
1 Introduction
[…]
In this current observational circumstance, a question arises; what kind of restrictions are generally placed on CP phases in the lepton sector, when $\delta$ is or is not observed by these experiments? For the CKM matrix, general analyses exist for hierarchical Yukawa matrices [5, 6] and many flavor texture studies [7]. However, no general treatment for each unitary matrix has been found for the two flavor mixing matrices $V_{\rm CKM}=U_{u}^{\dagger}U_{d}$ and $U_{\rm MNS}=U_{e}^{\dagger}U_{\nu}$. Therefore, in this letter, we analyze the quark and lepton mixing matrices in an almost general form for hierarchical mass matrices of the standard model fermions and discuss the distinguishability of CP phases of leptons by the next-generation experiments.
2 An almost general representation of flavor mixing matrices for hierarchical Yukawa structure
This section introduces an almost general form of flavor mixing matrices that comes from the hierarchical Yukawa matrices. At first, we define that a Yukawa matrix $Y_{f}$ of the Standard Model fermions $f=u,d,\nu,e$ is hierarchical if its elements satisfy $Y_{fii}\gg Y_{fij},Y_{ji}$ for $i>j$. Namely, for a matrix,
$Y_{f}\equiv\begin{pmatrix}Y_{f11}&Y_{f12}&Y_{f13}\\ Y_{f21}&Y_{f22}&Y_{f23}\\ Y_{f31}&Y_{f32}&Y_{f33}\\ \end{pmatrix},$ (1)
its elements $Y_{fij}$ satisfy
$Y_{f33}\gg Y_{f23},Y_{f32},Y_{f13},Y_{f31}\,,~{}Y_{f22}\gg Y_{f21},Y_{f12}\,.$ (2)
With these conditions, unitary matrices of the singular value decomposition of $Y_{f}$ necessarily has only small mixings, and singular values $m_{fi}$ of $Y_{f}$ are close to its diagonal elements $|Y_{fii}|\simeq m_{fi}$.
These hierarchical Yukawa matrices have approximate chiral flavor symmetries $U(1)_{L}^{2}\times U(1)_{R}^{2}$;
$Y_{f}^{\prime}=\begin{pmatrix}e^{i\alpha_{1L}}&0&0\\ 0&e^{i\alpha_{2L}}&0\\ 0&0&1\\ \end{pmatrix}Y_{f}\begin{pmatrix}e^{i\alpha_{1R}}&0&0\\ 0&e^{i\alpha_{2R}}&0\\ 0&0&1\\ \end{pmatrix}\simeq Y_{f}\,.$ (3)
Such $Y_{f}$ is easily realized by hierarchical Higgs vacuum expectation values (vevs) with appropriate chiral charges. Specifically, we consider a situation where the Yukawa matrices are generated by spontaneous symmetry breaking of flavon fields $\phi_{i}$ and $\tilde{\phi}_{j}$ that have chiral charges associated with the left-handed $i$ and right-handed $j$ generations. At the leading-order, Yukawa matrices are generated as
$Y_{f}={1\over\Lambda^{2}}\begin{pmatrix}y_{f11}\langle{\phi_{1}}\rangle\langle{\tilde{\phi}_{1}}\rangle&y_{f12}\langle{\phi_{1}}\rangle\langle{\tilde{\phi}_{2}}\rangle&y_{f13}\langle{\phi_{1}}\rangle\langle{\tilde{\phi}_{3}}\rangle\\ y_{f21}\langle{\phi_{2}}\rangle\langle{\tilde{\phi}_{1}}\rangle&y_{f22}\langle{\phi_{2}}\rangle\langle{\tilde{\phi}_{2}}\rangle&y_{f23}\langle{\phi_{2}}\rangle\langle{\tilde{\phi}_{3}}\rangle\\ y_{f31}\langle{\phi_{3}}\rangle\langle{\tilde{\phi}_{1}}\rangle&y_{f32}\langle{\phi_{3}}\rangle\langle{\tilde{\phi}_{2}}\rangle&y_{f33}\langle{\phi_{3}}\rangle\langle{\tilde{\phi}_{3}}\rangle\\ \end{pmatrix}.$ (4)
Here, $y_{fij}$ denotes coupling constants between flavons and fermions $f$, and $\Lambda$ is some cutoff scale. Approximate chiral symmetries exist if these vevs satisfy $\langle{\phi_{3}}\rangle\gg\langle{\phi_{2}}\rangle\gg\langle{\phi_{1}}\rangle$ and $\langle{\tilde{\phi}_{3}}\rangle\gg\langle{\tilde{\phi}_{2}}\rangle\gg\langle{\tilde{\phi}_{1}}\rangle$. Various mass matrix models exhibit these properties.
2.1 An assumption from hierarchy of Yukawa matrices
For the aforementioned hierarchical Yukawa matrices, this letter deals with two flavor mixing matrices, the CKM matrix $V_{\rm CKM}$ and the MNS matrix $U_{\rm MNS}$.
$V_{\rm CKM}=U_{u}^{\dagger}U_{d}\,,~{}~{}~{}U_{\rm MNS}=U_{e}^{\dagger}U_{\nu}\,.$ (5)
Here, $U_{f}$ denote left-handed unitary matrices that diagonalize $Y_{f}$ and the neutrino mass matrix $m_{\nu}$. This $U_{f}$ is generally displayed in the PDG parameterization and its mixing angle $s_{ij}^{f}$.
To make the analysis concise, the following assumption is imposed.
Assumption:
We neglect the 1-3 mixing of $U_{f}$ for fermions $f$ with a stronger hierarchy of singular values (i.e. $f=u,e$),
This is justified by the approximate chiral symmetries described above, because the 1-3 mixing of $U_{u,e}$ is suppressed by $m_{f1}/m_{f2},m_{f1}/m_{f3}$.
To be specific, if the unitary matrices diagonalizing $Y_{f}$ have small mixings, approximate chiral symmetries roughly constrain the texture of $Y_{f}$ as
$Y_{f}\simeq(Y_{f})_{33}\begin{pmatrix}\epsilon_{L}^{f}\epsilon_{R}^{f}&\epsilon_{L}^{f}\delta_{R}^{f}&\epsilon_{L}^{f}\\ \delta_{L}^{f}\epsilon_{R}^{f}&\delta_{L}^{f}\delta_{R}^{f}&\delta_{L}^{f}\\ \epsilon_{R}^{f}&\delta_{R}^{f}&1\end{pmatrix},$ (6)
where $\epsilon_{L,R}^{f},\delta_{L,R}^{f}$ are breaking parameters of the $U(1)_{(L,R)}$ symmetry for the first and second generations, and $O(1)$ coefficients are omitted. The behavior of the 1-3 mixing is divided into three cases depending on the magnitude of $\epsilon_{L,R}^{f}$.
  1. 1.
    $|\epsilon_{R}^{f}|\sim 1\gg|\epsilon_{L}^{f}|$: The 1-3 mixings are approximately evaluated as
    $|s_{13}^{e}|$ $\simeq|\epsilon_{L}|\simeq m_{e1}/m_{e3}\simeq 0.0003\,,$ (7)
    $|s_{13}^{u}|$ $\simeq|\epsilon_{L}|\simeq m_{u1}/m_{u3}\simeq 0.00001\,.$ (8)
    We can neglect them because these mixings are sufficiently smaller than the 1-3 elements of the flavor mixing matrices $|(V_{\rm CKM})_{13}|\sim 10^{-3},|(U_{\rm MNS})_{13}|\sim 10^{2}$. Besides, for too large $\epsilon_{R}^{f}\gtrsim 0.1$, there is no longer an approximate symmetry.
  2. 2.
    $|\epsilon_{L}^{f}|\sim|\epsilon_{R}^{f}|$: Similarly,
    $|s_{13}^{e}|$ $\simeq|\epsilon_{L}|\simeq\sqrt{m_{e1}/m_{e3}}\simeq 0.017\,,$ (9)
    $|s_{13}^{u}|$ $\simeq|\epsilon_{L}|\simeq\sqrt{m_{u1}/m_{u3}}\simeq 0.0027\,.$ (10)
    Therefore, its contribution for leptons is at most a $\pm$10change in $s_{13}$. On the other hand, for quarks, the 1-3 mixing can have a large impact on $|V_{ub}|\simeq 0.003$. However, a similar hierarchy for down-type quarks leads to $\sqrt{m_{d1}/m_{d3}}\simeq 0.03$, and such a model with large 1-3 mixings is excluded unless a fine-tuned cancellation.
  3. 3.
    $|\epsilon_{L}|\gg|\epsilon_{R}|$: In this case, it is also hard to consider that there is an approximate $U(1)_{L}$ symmetry. The assumption does not hold and the unitary matrix for left-handed fields has large mixings. With the corresponding CKM matrix element $|V_{ub}|\simeq 0.003$ in mind, this situation seems difficult to realize in a unified theory.
Thus, this approximation is justified in a wide range of parameters.
The assumption eliminates the intrinsic CP violation associated with $U_{u,e}$, because three generations are no longer involved in flavor mixing for the more hierarchical fermions. Thus, in this analysis, there are two sources of CP violation: relative phases between the two unitary matrices and intrinsic CP phase for the non-hierarchical fermions.
A similar discussion allows some mention for magnitudes of the 1-2 mixings $\epsilon_{L}^{f}/\delta_{L}^{f}$.
${m_{e1}\over m_{e2}}\simeq 0.005\lesssim|s_{12}^{e}|\lesssim 0.07\simeq\sqrt{m_{e1}\over m_{e2}}\,,$ (11)
${m_{u1}\over m_{u2}}\simeq 0.002\lesssim|s_{12}^{u}|\lesssim 0.04\simeq\sqrt{m_{u1}\over m_{u2}}\,.$ (12)
In this letter, we will proceed with the analysis by considering $s_{12}^{e,u}\lesssim 0.1$.
2.2 Representation of flavor mixing matrices
Using this assumption, we organize the representation of the MNS (and CKM) matrix. For $U_{\nu,e}$ displayed in the PDG parameterization, the determinant of 2-3 submatrix can be made to unity by a redefinition of the overall phase. Then the MNS matrix is written by
$U_{\rm MNS}$ $\equiv O_{e}^{T}\begin{pmatrix}e^{i\rho}&0&0\\ 0&e^{i\sigma}&0\\ 0&0&e^{-i\sigma}\\ \end{pmatrix}O_{\nu}P,~{}~{}~{}O_{e}^{T}\equiv\begin{pmatrix}c_{e}&-s_{e}&0\\ s_{e}&c_{e}&0\\ 0&0&1\\ \end{pmatrix}\begin{pmatrix}1&0&0\\ 0&c_{\tau}&-s_{\tau}\\ 0&s_{\tau}&c_{\tau}\\ \end{pmatrix}\,,$
$O_{\nu}P$ $\equiv\begin{pmatrix}1&0&0\\ 0&c_{\nu}&s_{\nu}\\ 0&-s_{\nu}&c_{\nu}\\ \end{pmatrix}\begin{pmatrix}c_{13}&0&s_{13}e^{-i\delta_{\nu}}\\ 0&1&0\\ -s_{13}e^{i\delta_{\nu}}&0&c_{13}\\ \end{pmatrix}\begin{pmatrix}c_{12}&s_{12}&0\\ -s_{12}&c_{12}&0\\ 0&0&1\\ \end{pmatrix}\begin{pmatrix}e^{i\phi_{1}}&0&0\\ 0&e^{i\phi_{2}}&0\\ 0&0&e^{i\phi_{3}}\\ \end{pmatrix}.$ (13)
Here $s_{f}\equiv\sin f\,,\,c_{f}\equiv\cos f\,,\,s_{ij}\equiv\sin\theta_{ij}\,,\,c_{ij}\equiv\cos\theta_{ij}$ and $\rho,\sigma,\delta_{\nu},\phi_{i}$ are arbitrary phases. Since the phases $\phi_{i}$ only affect the Majorana phases, they are omitted hereafter.
The two 2-3 mixings are merged to define new arguments $\chi$ and $\eta$.
$\begin{pmatrix}c_{\tau}&-s_{\tau}\\ s_{\tau}&c_{\tau}\\ \end{pmatrix}\begin{pmatrix}e^{i\sigma}&0\\ 0&e^{-i\sigma}\\ \end{pmatrix}\begin{pmatrix}c_{\nu}&s_{\nu}\\ -s_{\nu}&c_{\nu}\\ \end{pmatrix}=\begin{pmatrix}e^{i\sigma}c_{\tau}c_{\nu}+e^{-i\sigma}s_{\tau}s_{\nu}&e^{i\sigma}c_{\tau}s_{\nu}-e^{-i\sigma}s_{\tau}c_{\nu}\\ e^{i\sigma}s_{\tau}c_{\nu}-e^{-i\sigma}c_{\tau}s_{\nu}&e^{i\sigma}s_{\tau}s_{\nu}+e^{-i\sigma}c_{\tau}c_{\nu}\\ \end{pmatrix}$ (14)
$=\begin{pmatrix}c_{\sigma}c_{\nu-\tau}+is_{\sigma}c_{\nu+\tau}&c_{\sigma}s_{\nu-\tau}+is_{\sigma}s_{\nu+\tau}\\ -c_{\sigma}s_{\nu-\tau}+is_{\sigma}s_{\nu+\tau}&c_{\sigma}c_{\nu-\tau}-is_{\sigma}c_{\nu+\tau}\\ \end{pmatrix}\equiv\begin{pmatrix}e^{i\chi}c_{23}&e^{i\eta}s_{23}\\ -e^{-i\eta}s_{23}&e^{-i\chi}c_{23}\\ \end{pmatrix}\,.$ (15)
Relations between these arguments $\chi,\eta$ and the phases are
$\tan\chi=\tan\sigma{c_{\nu+\tau}\over c_{\nu-\tau}}\,,~{}~{}~{}\tan\eta=\tan\sigma{s_{\nu+\tau}\over s_{\nu-\tau}}\,.$ (16)
From this,
$U_{\rm MNS}$ $=\begin{pmatrix}c_{e}&-s_{e}&0\\ s_{e}&c_{e}&0\\ 0&0&1\\ \end{pmatrix}\begin{pmatrix}e^{i\rho}&0&0\\ 0&e^{i\chi}c_{23}&e^{i\eta}s_{23}\\ 0&-e^{-i\eta}s_{23}&e^{-i\chi}c_{23}\\ \end{pmatrix}\begin{pmatrix}c_{13}&0&s_{13}e^{-i\delta_{\nu}}\\ 0&1&0\\ -s_{13}e^{i\delta_{\nu}}&0&c_{13}\\ \end{pmatrix}\begin{pmatrix}c_{12}&s_{12}&0\\ -s_{12}&c_{12}&0\\ 0&0&1\\ \end{pmatrix}\,,$ (17)
Furthermore, by multiplying diag $(1,1,e^{i(\chi+\eta)})$ from the left and diag $(1,1,e^{i(\chi-\eta)})$ from the right of $U_{\rm MNS}$ and removing an overall phase $e^{i\chi}$, we obtain
$U_{\rm MNS}$ $=\begin{pmatrix}c_{e}&-s_{e}&0\\ s_{e}&c_{e}&0\\ 0&0&1\\ \end{pmatrix}\begin{pmatrix}e^{i(\rho-\chi)}&0&0\\ 0&c_{23}&s_{23}\\ 0&-s_{23}&c_{23}\\ \end{pmatrix}\begin{pmatrix}c_{13}&0&s_{13}e^{-i(\delta_{\nu}+\eta-\chi)}\\ 0&1&0\\ -s_{13}e^{i(\delta_{\nu}+\eta-\chi)}&0&c_{13}\\ \end{pmatrix}\begin{pmatrix}c_{12}&s_{12}&0\\ -s_{12}&c_{12}&0\\ 0&0&1\\ \end{pmatrix}\,.$ (18)
Therefore, the following two phases $\alpha$ and $\beta$ are physical.
$\alpha\equiv\rho-\chi\,,~{}~{}~{}\beta=\delta_{\nu}+\eta-\chi\,.$ (19)
If $s_{\tau}$ is sufficiently small by the chiral symmetry, $\chi\sim\eta\sim\sigma$ and $\alpha\sim\rho-\sigma\,,\,\beta\sim\delta_{\nu}$ hold. This notation is a generalization of the Fritzsch–Xing parameterization [8].
Finally, we discuss sign degrees of freedom for mixing angles. Continuous parameters in this notation are $s_{e},s_{ij}$ and two phases. Similar to the PDG parameterization, phase redefinitions of the six leptons can eliminate five signs. Since the sign of $s_{13}$ can be absorbed into the phase $e^{\pm i\beta}$, the six $s_{ij}$ and $c_{ij}$ are set to be positive. By multiplying diag $(-1,-1,1)$ from the left, the sign of $c_{e}$ is inverted without changing the sign of $s_{ij}$ and $c_{ij}$. Therefore, only the sign of $s_{e}$ is physical because the sign of the two $s_{e}$ cannot be absorbed into the phase $\alpha$. We can specifically check that the physical sign in the CP violation is sign$(c_{e}s_{e})$ later. Next subsection shows that there are three free parameters $\alpha,\beta,s_{e}$ in the MNS matrix, and two in the CKM matrix because the CP phase is further determined.
2.3 Constraints on mixing angles
We will consider how the observables constrain these parameters. The best-fit values of the latest global fit [9] are used as input parameters for the MNS matrix. The three mixing angles of the normal hierarchy (NH) without Super-Kamiokande (SK) are
$\sin^{2}\theta_{12}^{\rm NH}=0.304\,,~{}~{}\sin^{2}\theta_{23}^{\rm NH}=0.573\,,~{}~{}~{}\sin^{2}\theta_{13}^{\rm NH}=0.0222\,.$ (20)
The reason for using NH without SK is that the values of inverted hierarchy (IH) with or without SK are close to these values. Although the inclusion of the SK data makes $s_{23}^{\rm PDG}$ about $0.1$ smaller for NH, the qualitative behavior in the following discussion remains the same.
Three mixing parameters $s_{ij}$ are constrained from these input parameters. Hereafter, the above three input parameters displayed in PDG parameterization are denoted as $\sin\theta_{ij}^{\rm PDG}\equiv S_{ij}\,,\cos\theta_{ij}^{\rm PDG}\equiv C_{ij}$. First, the 3-3 element of $U_{\rm MNS}$, which is the same as the PDG parameterization, determines $s_{23}$;
$c_{13}c_{23}=C_{13}C_{23}\,,~{}~{}~{}s_{23}=\sqrt{1-{C_{13}^{2}C_{23}^{2}\over c_{13}^{2}}}\,.$ (21)
Next, $|U_{13}|^{2}$ is a function of $s_{13}$ and $s_{23}$.
$U_{13}$ $=c_{e}s_{13}e^{i(\alpha-\beta)}-c_{13}s_{23}s_{e}\,,$ (22)
$|U_{13}|^{2}$ $=c_{e}^{2}s_{13}^{2}+c_{13}^{2}s_{23}^{2}s_{e}^{2}-2c_{13}c_{e}s_{13}s_{23}s_{e}\cos(\alpha-\beta)\,.$ (23)
Substituting the solution of $s_{23}$ and setting $|U_{13}|^{2}$ equal to $S_{13}^{2}$, we obtain a quadratic equation for $s_{13}^{2}$.
$s_{13}^{4}(4c_{e}^{2}s_{e}^{2}\cos^{2}(\alpha-\beta)+(s_{e}^{2}-c_{e}^{2})^{2})+(s_{e}^{2}(C_{13}^{2}C_{23}^{2}-1)+S_{13}^{2})^{2}$
$+s_{13}^{2}[2(s_{e}^{2}-c_{e}^{2})(s_{e}^{2}(C_{13}^{2}C_{23}^{2}-1)+S_{13}^{2})+4c_{e}^{2}s_{e}^{2}(C_{13}^{2}C_{23}^{2}-1)\cos^{2}(\alpha-\beta)]=0\,.$ (24)
The two solutions of $s_{13}^{2}$ are
$s_{13}^{2}$ $={A\pm\sqrt{B}\over(c_{e}^{4}+2c_{e}^{2}s_{e}^{2}\cos 2(\alpha-\beta)+s_{e}^{4})}\,,$ (25)
$A$ $=-c_{e}^{2}s_{e}^{2}(C_{13}^{2}C_{23}^{2}-1)\cos 2(\alpha-\beta)-C_{13}^{2}C_{23}^{2}s_{e}^{4}+c_{e}^{2}S_{13}^{2}-S_{13}^{2}s_{e}^{2}+s_{e}^{4}\,,$ (26)
$B$ $=2c_{e}s_{e}\cos(\alpha-\beta)\sqrt{C_{13}^{2}S_{13}^{2}S_{23}^{2}-c_{e}^{2}s_{e}^{2}(C_{13}^{2}C_{23}^{2}-1)^{2}\sin^{2}(\alpha-\beta)}\,.$ (27)
For sufficiently small $s_{13}$, since the two independent solutions are absorbed to the sign of $\cos(\alpha-\beta)$, the sign $A+\sqrt{B}$ is chosen as the solution that reproduces the correct $U_{13}$. In this case, the other solution corresponds to the dashed line in Fig. 1.
From the triangular inequality for Eq. (23), upper and lower bounds exist on $s_{13}$.
$c_{e}S_{13}-C_{13}S_{23}s_{e}\leqq s_{13}\leqq C_{13}S_{23}s_{e}+c_{e}S_{13}\,.$ (28)
This result is easy to understand because the mixing matrix of $s_{ij}$ is a “reverse rotation” of the PDG representation of $S_{ij}$. From this, there exists a critical value $s_{e}^{c}$ of $s_{e}$ such that $s_{13}=0$.
$s_{13}=0~{}~{}\Rightarrow~{}~{}{s_{e}^{c}\over c_{e}^{c}}={S_{13}\over C_{13}S_{23}}\,.$ (29)
This value is about 0.2 for the MNS matrix (and 0.09 for the CKM matrix). For larger $s_{e}>s_{e}^{c}$, as can be seen in Fig. 1, the two solutions in Eq. (25) (with different signs before the root sign) are physically inequivalent.
Figure 1: Geometric explanation of solutions of $s_{13}$ for given mixing angles $s_{23}$ and $s_{e}$. If $s_{e}$ is below the critical value, there are two solutions with different signs of $\cos(\alpha-\beta)$. If $s_{e}$ is above the critical value, there are two solutions with different magnitudes for the same $\cos(\alpha-\beta)$.
Finally, we consider the condition on $s_{12}$. The matrix element $U_{12}$ is explicitly written as
$U_{12}=-c_{12}c_{23}s_{e}+e^{i\alpha}c_{13}c_{e}s_{12}+e^{i\beta}s_{12}s_{23}s_{13}s_{e}\,.$ (30)
By substituting the solutions of $s_{23}$ (21) and $s_{13}$ (25) to the constraint $|U_{12}|^{2}=C_{13}^{2}S_{12}^{2}$, $s_{12}^{2}$ is similarly obtained by solving a quadratic equation. As in the case of $s_{13}$, the sign before the root sign can be absorbed into the term $2c_{23}s_{e}(c_{13}c_{e}\cos\alpha+s_{13}s_{23}s_{e}\cos\beta)$.
In particular, the dependence of $\beta$ on $s_{12}$ is small, because it is suppressed by $s_{13}s_{e}$. Thus, the constraint on $s_{12}$ is approximately
$C_{13}^{2}S_{12}^{2}\simeq(c_{12}c_{23}s_{e})^{2}+(c_{13}c_{e}s_{12})^{2}-2c_{12}c_{23}s_{e}c_{13}c_{e}s_{12}\cos\alpha\,.$ (31)
Since this constraint has a similar structure to Eq. (23), we choose $s_{12}\simeq A^{\prime}+\cos\alpha\,B^{\prime}>0$ as a positive solution that reproduces the matrix elements of $U_{\rm MNS}$. Here, $A^{\prime},B^{\prime}$ are functions of several trigonometric functions. We will not discuss the critical value of $s_{12}$ because $s_{12}$ is relatively large for both the CKM and MNS matrices.
By imposing the three constraints, we can treat CP violation as a function of $\alpha,\beta$ and $s_{e}$. The rough parameter dependences of the three solutions are as follows.
  • $s_{23}$: From Eq. (21), dependences of $\alpha,\beta$ on $s_{23}$ are relatively small because these phases enter in the second order of $s_{e}$ and $s_{13}$.
    $s_{23}=S_{23}+O(s_{13}^{2},s_{e}s_{13},s_{e}^{2})\,.$ (32)
  • $s_{13}$: From Eq. (23), $s_{13}$ depends on $\alpha-\beta$. The expansion for small $s_{e}$ is
    $s_{13}=S_{13}+C_{13}S_{23}s_{e}\cos(\alpha-\beta)+O(s_{e}^{2})\,.$ (33)
    However, $s_{e}/s_{13}$ is often not negligible, and this expansion does not have good accuracy.
  • $s_{12}$: Since $\beta$ dependence is suppressed by $s_{13}s_{e}$ in Eq. (30), $s_{12}$ has an almost only $\alpha$ dependence. Therefore, the behavior of the solution is
    $s_{12}\simeq S_{12}+C_{12}C_{23}s_{e}\cos\alpha\,.$ (34)
3 The Jarlskog invariant and its numerical results
In this section, we investigate the behavior and numerical results of CP violation after imposing the three experimental constraints. The Jarlskog invariant $J$ [10] of the unitary matrix $U_{\rm MNS}$ (18) is
$J$ $={\rm Im\,}[U_{\mu 2}U_{\tau 3}U_{\mu 3}^{*}U_{\tau 2}^{*}]$ (35)
$=c_{13}c_{23}[c_{12}s_{12}\left(c_{13}s_{23}s_{13}(c_{e}^{2}-s_{e}^{2})\sin\beta-c_{13}^{2}s_{23}^{2}c_{e}s_{e}\sin\alpha\right)$
$+c_{23}(c_{12}^{2}-s_{12}^{2})s_{23}s_{13}c_{e}s_{e}\sin(\alpha-\beta)+s_{13}^{2}c_{e}s_{e}\left(c_{23}^{2}\sin\alpha-s_{23}^{2}\sin(\alpha-2\beta)\right)]\,.$ (36)
To the extent that $s_{e},s_{13}\lesssim 0.1$ holds, higher orders of $s_{e}$ and $s_{13}$ can be negligible, and the dominant terms are those proportional to $\sin\alpha$ and $\sin\beta$. Note that the relative sign between terms depends only on $s_{e}c_{e}$.
From Eqs. (32), (33), and (34), behavour of the Dirac phase $\delta$ for small $s_{e}$ is
$\sin\delta$ $={J\over C_{12}S_{12}C_{23}S_{23}C_{13}^{2}S_{13}}$ (37)
$\simeq\sin\beta+s_{e}\left(-{S_{23}\cos\beta\sin(\alpha-\beta)\over S_{13}}+{(1-2S_{12}^{2})C_{23}\cos\alpha\sin\beta\over C_{12}S_{12}}\right).$ (38)
The term with $\sin(\alpha-\beta)$ is a sum of the term with $\sin\alpha$ and a contribution from the expansion of $s_{13}$, while the term with $\cos\alpha$ is from the expansion of $s_{12}$. In the MNS matrix, the second term is dominant over the third term because $S_{23}/S_{13}\simeq 5$ and $(1-2S_{12}^{2})C_{23}/C_{12}S_{12}\simeq 1/2$ hold. The same is true for the CKM matrix from $S_{23}/S_{13}\simeq 10$.
3.1 Numerical results
Plots of $\sin\delta$ as a function of $\alpha,\beta$ and $s_{e}$ are shown in Fig. 2. The range of $s_{e}$ is set within the critical value $|s_{e}|\leqq 0.2$. The green-shaded region corresponds to $\cos\delta<0$. From Eq. (38), if $s_{e}$ is sufficiently small, $\delta$ is approximately equal to $\beta$. On the other hand, when $s_{e}$ approaches the critical value, it shows singular behavour on the line $\alpha-\beta=\pm\pi/2,\pm 3\pi/2$. For $s_{e}$ above the critical value, the divided two regions correspond to the large and the small solution, respectively.
Figure 2: Plots of $\sin\delta$ of the MNS matrix $U_{\rm MNS}$ for a given 1-2 mixing $s_{e}$ and phases $\alpha,\beta$. The green-shaded regions represent $\cos\delta<0$.
In particular, for $\beta=0$ or $\pi$, a kind of generalized CP symmetry (GCP) [11, 12, 13, 14, 15] can be defined. In this case, there is an upper bound on $\delta$ from Eq. (38).
$|\sin\delta^{\rm max}|\simeq\left|{s_{e}S_{23}\over S_{13}}\sin\alpha\right|\,.$ (39)
However, this expression is not accurate for $s_{e}\simeq s_{13}$, and a large CPV such as $\delta\simeq 1$ is possible even for $\beta=0$ or $\pi$ and $s_{e}\sim s_{e}^{c}$. On the other hand, the chiral symmetry does not hold for $s_{e}\simeq 0.2$, and the parameter region where $s_{e}$ is close to the critical value is out of scope in this letter.
The long-baseline experiment using HK and J-PARC [3] has a discovery capability of around $\delta\gtrsim 30^{\circ}$ after 10 years of operation. This allows for physical interpretations when HK observes or does not observe the Dirac phase $\delta$, respectively.
When HK observes the Dirac phase (of about $O(1)$)
Large CP violation, as suggested by the T2K experiment, cannot be achieved by the relative phase $\alpha$ alone. Therefore, if the Yukawa matrix of charged leptons $Y_{e}$ is sufficiently hierarchical, large CP violation corresponds to the intrinsic CP violation $\delta_{\nu}$ in the diagonalization $O_{\nu}$ of the neutrino matrix. In particular, if the phase is close to $\pi/2$, $\mu-\tau$ reflection symmetry [16, 17, 18] (and as a review [19]) is more favored.
When HK does not observe the Dirac phase (of about 30 or more)
[…]
${s_{e}S_{23}\over S_{13}}\sin\alpha\,,~{}\sin\beta<\sin\delta^{\rm lim}\sim 0.5\,.$ (40)
In such a situation, a kind of CP symmetry such as $\beta=0$ or $\pi$ can be favored. However, even for large $s_{e}\sim 0.1$, the relative phase $\sin\alpha$ is hardly restricted (and $\alpha\sim\pm\pi/2$ is allowed). The smaller the observed CPV, the more important the relative phase can be for the source of CPV. Such restrictions on the CP phase can provide important constraints on the leptogenesis scenario [20].
3.2 CKM matrix
A similar analysis is performed for the CKM matrix. As input parameters, we use the latest data of UTfit [21].
$\sin\theta_{12}$ $=0.22519\pm 0.00083\,,~{}~{}~{}\sin\theta_{23}=0.04200\pm 0.00047\,,$ (41)
$\sin\theta_{13}$ $=0.003714\pm 0.000092\,,~{}~{}~{}\delta_{\rm KM}=1.137\pm 0.022\,.$ (42)
The parameter $s_{e}$ in the MNS matrix is redefined as $s_{u}$ in this case, and a similar analysis for the CKM matrix is shown in Fig. 3. The range of $s_{u}$ is set to be within the critical value ($|s_{u}|\leqq 0.09$) and the region within the error range of $\delta_{\rm KM}$ was filled in. The green-shaded regions are excluded because $\cos\delta_{\rm KM}<0$. In the three continuous parameters, one two-dimensional surface is chosen that reproduces the observed CP phase.
Figure 3: Plots of the CP violation $\sin\delta_{\rm KM}$ of the quark mixing matrix $V_{\rm CKM}$. The green-shaded regions are excluded because they represent $\cos\delta_{\rm KM}<0$.
Since the phase behaves as $\delta_{\rm KM}\simeq\beta$ for sufficiently small $s_{u}$, we already see that the observed CPV in the CKM matrix corresponds to the intrinsic CP phase associated with the diagonalization of down quarks. On the other hand, near the critical value of $s_{u}\simeq 0.09$, a parameter set $\beta=0$ or $\pi$ and $\alpha\simeq\pm\pi/2$ is possible. In the limit of $s_{13}\to 0$, Eq. (18) is reduced to the Fritzsch–Xing parameterization and it has been known that the phase is almost maximal $\alpha=\pi/2$ [8]. This situation is realized by a kind of generalized CP symmetries [22, 23, 24, 25]. Such restrictions can be useful in the construction of grand unified theories.
4 Summary
In this letter, we perform an almost general analysis of flavor mixing matrices to investigate the discriminative power of CP phases by the next-generation neutrino oscillation experiments. For hierarchical Yukawa matrice $Y_{f}$ of the Standard Model fermions $f$, the unitary matrix $U_{f}$ of diagonalization for left-handed fields is expected to have small mixings. As an assumption, we neglect the 1-3 mixing of more hierarchical fermions (i.e. up-type quarks and charged leptons). Thus there are two sources of CP violation in the flavor mixing matrices, the intrinsic CP phase $\delta_{d,\nu}$ in diagonalization of less hierarchical fermions $d,\nu$ and relative phases between two unitary matrices.
By eliminating unphysical phases and imposing constraints of the three measured mixing angles, the flavor mixing matrices are analytically displayed by two phases and the 1-2 mixing $s_{u,e}$ of more hierarchical fermions. For sufficiently small 1-2 mixing $s_{e}$ of charged leptons, the Dirac phase $\delta$ is almost identical to the intrinsic phase of neutrinos $\delta_{\nu}$ and contributions of relative phases are not dominant. Therefore, future detection of the Dirac phase by the Hyper-Kamiokande experiment indicates the observation of $\delta_{\nu}$. On the other hand, if such a CP violation is not observed, an upper limit is placed on a combination of $\delta_{\nu}$ and relative phases.
The analysis is also performed for the CKM matrix and similar conclusions are obtained. Such restrictions will be useful in analyses of leptogenesis and grand unified theories.